# Solving the Hubbard model with Neural Quantum States

### v1

@@ 2 Pretrain
Instead of randomly initializing the NQS ansatz, we use a pretrain method to provide the transformer wavefuntion a better initial parameters before optimizing it with MARCH.
We first train a simple neural network similar to neural network backflow (NNB)  (?). A multilayer perceptron (MLP) with two hidden layers maps the occupation number $\mathbf{n}\in\{0,1\}^{2N}$ to a orbital matrix $\mathbf{M}_{\text{nnb}}(\mathbf{n})\in\mathbb{R}^{2N\times N_{e}}$, which is restricted to a $2\times 2$ block-diagonal matrix,
$\mathbf{M}_{\text{nnb}}(\mathbf{n})=\begin{pmatrix}\mathbf{M}_{\text{nnb}}^{\uparrow}(\mathbf{n})&0\\ 0&\mathbf{M}_{\text{nnb}}^{\downarrow}(\mathbf{n})\end{pmatrix},$ (S24)
where the $\mathbf{M}_{\text{nnb}}^{\uparrow}(\mathbf{n}),\mathbf{M}_{\text{nnb}}^{\downarrow}(\mathbf{n})\in\mathbb{R}^{N\times N_{e}/2}$ are the orbitals of spin-up and spin-down electrons, respectively. Based on the occupation number, $N_{e}$ rows of orbital matrix $\mathbf{M}_{\text{nnb}}(\mathbf{n})$ are extracted to form a square orbital matrix $\mathbf{\Phi}\in\mathbb{R}^{N_{e}\times N_{e}}$, which is also $2\times 2$ block-diagonal,
$\mathbf{\Phi}=\begin{pmatrix}\mathbf{\Phi}^{\uparrow}&0\\ 0&\mathbf{\Phi}^{\downarrow}\end{pmatrix}.$ (S25)
The amplitude of our Neural Network-Based wavefunction ansatz is defined by the determinant of a matrix $\mathbf{\Phi}$:
$\psi_{\text{nnb}}(\mathbf{n})=\mathrm{det}[\mathbf{\Phi}].$ (S26)
@@
$S_{\text{pre}}=\left\{\mathbf{n}_{i},\ \mathbf{M}_{\text{nnb}}(\mathbf{n}_{i})\right\}_{i=0}^{N_{\text{pre}}},$ (S27)
where each entry consists of a configuration $\mathbf{n}_{i}$ and its corresponding orbital matrix $\mathbf{M}_{\text{nnb}}(\mathbf{n}_{i})$.
Transformer ansatz is then pretrained in a supervised manner, with the loss function to be the mean square error of the orbital matrix, i.e.,
$\mathcal{L}_{\text{pretrain}}=\frac{1}{N_{\text{pre}}}\sum_{i=1}^{N_{\text{pre}}}||\mathbf{M}_{\text{Transformer}}(\mathbf{n}_{i})-\mathbf{M}_{\text{nnb}}(\mathbf{n}_{i})||^{2},$ (S28)
with Adam optimizer. We note that in practice, the pretrain dataset can be generated online. Compared to random initialization, pretrained transformer is much more stable and converges significantly faster.
@@ 3 Pinning Field
To avoid convergence to local minima during optimization, a temporary pinning field is applied. This field takes the form of an antiferromagnetic (AFM) term added to the Hamiltonian on specific columns:
@@
4 Green Function Monte Carlo
The Green Function Monte Carlo (GFMC) aims to project out the ground state wavefunction $\psi_{0}$ from a trial wavefunction $\psi_{T}$, by iteratively applying the imaginary time evolution operator $e^{-\tau(\hat{H}-E)}$, where $E$ is an energy offset and $\tau$ is a small time step. In practice, this evolution is often implemented using first-order approximation $T=1-\tau(\hat{H}-E)$. A significant challenge arises because the Hamiltonian can have negative off-diagonal matrix elements, leading to negative values in the projector $T$. This ”sign problem” prevents a direct interpretation of the wavefunction amplitude as a probability distribution, rendering standard Monte Carlo sampling exponentially hard.
Following  (?, ?), we apply fixed node approximation to enforce positivity, albeit at the cost of introducing a systematic error unless the trial wavefunction is exact. The core idea is to construct an effective Hamiltonian $\hat{H}_{\text{eff}}$. We first identify problematic transitions, termed Sign-Flips (SF), between configuration $R$ and $R^{\prime}$ with $R\neq R^{\prime}$ if the following condition holds:
$\bra{R}\hat{H}\ket{R^{\prime}}\psi_{T}(R)\psi_{T}(R^{\prime})>0$ (S30)
This condition identifies off-diagonal elements that, when combined with the signs of the trial wavefunction at the respective configurations, is problematic for the specific GFMC formulation. The effective Hamiltonian $\hat{H}_{\text{eff}}$ is then constructed as follows:
[…]
Essentially, off-diagonal elements corresponding to SF transition are removed. To compensate, the diagonal elements $\bra{R}\hat{H}_{\text{eff}}\ket{R}$ are adjusted by adding terms derived from the removed SF transitions.
To improve sampling efficiency, importance sampling using the trial wavefunction $\psi_{T}$ is employed. Instead of directly simulating the evolution governed by $T=1-\tau(\hat{H}_{\text{eff}}-E)$, we consider the evolution of the distribution $f(R,t)=\psi(R,t)\psi_{T}(R)$. The corresponding importance-sampled transition matrix $\tilde{T}(R^{\prime},R)$ is defined as $\tilde{T}(R^{\prime},R)=T(R^{\prime},R)\frac{\psi_{T}(R^{\prime})}{\psi_{T}(R)}$. The explicit form of this matrix is:
$\tilde{T}(R^{\prime},R)=\begin{cases}1-\tau(\bra{R}\hat{H}_{\text{eff}}\ket{R^{\prime}})+\tau E,\quad\text{if }R^{\prime}=R\\ 0,\quad\text{if }(R,R^{\prime})\text{ is SF}\\ -\tau\bra{R^{\prime}}\hat{H}\ket{R}\frac{\psi_{T}(R^{\prime})}{\psi_{T}(R)}\quad\text{else}\end{cases}$ (S32)
Once the evolution has converged, the expectation value of any operator $\hat{A}$ can be evaluated. If $\hat{A}$ commutes with the Hamiltonian $\hat{H}$, its expectation value is calculated using the mixed estimator:
$\braket{\hat{A}}_{\text{mixed}}=\frac{\bra{\psi_{0}}\hat{A}\ket{\psi_{T}}}{\braket{\psi_{0}}{\psi_{T}}}=E_{R\sim\psi_{0}\psi_{T}}\left[\frac{\hat{A}\psi_{T}(R)}{\psi_{T}(R)}\right].$ (S33)
Specifically, when $\hat{A}$ is the Hamiltonian $\hat{H}$, the Fixed-Node approximation guarantees that the energy obtained from the effective Hamiltonian $\hat{H}_{\text{eff}}$ is an upper bound to the true ground state energy $E_{0}$, as proven in  (?). Consequently, by utilizing a variationally optimized NQS as the trial wavefunction, the GFMC method can yield a superior variational energy.
Also in principle, exact pure estimators of local operators can be obtained via the future-walking algorithm  (?). The core idea is to rewrite the pure expectation value as
[…]
where $\frac{\psi_{0}(R)}{\psi_{T}(R)}$ can be obtained from the asymptotic offspring of the walker:
$\frac{\psi_{0}(R)}{\psi_{T}(R)}$ $=\lim_{t\to\infty}\bra{R}T^{t}\ket{\psi_{T}}/\psi_{T}(R)$ (S35)
$=\lim_{t\to\infty}\sum_{R_{1},...R_{t}}T(R_{1},R)T(R_{2},R_{1})\cdots T(R_{t},R_{t-1})\frac{\psi_{T}(R_{t})}{\psi_{T}(R)}$ (S36)
$=\lim_{t\to\infty}\sum_{R_{1},...R_{t}}\tilde{T}(R_{1},R)\tilde{T}(R_{2},R_{1})\cdots\tilde{T}(R_{t},R_{t-1})$ (S37)
$=W(R)$ (S38)
where $W(R)$ is the sum of the weight of all descendants of $R$, and can be estimated directly by Monte Carlo sampling. Therefore,
$\braket{A}_{\text{pure}}=E_{R\sim\psi_{0}\psi_{T}}\frac{A\psi_{T}(R)W(R)}{\psi_{T}(R)W(R)}$ (S39)
@@ 5.2 Comparison with DMRG
We further compare our NQS method with the DMRG on the $t^{\prime}$ Hubbard model. Specifically, for a cylindrical geometry $16\times 8$ with next-nearest-neighbor hopping $t^{\prime}=-0.2$, our NQS achieves a variational energy of $-0.72791$. This is lower than the DMRG energy of $-0.72747$, obtained with a large bond dimension of $m=40000$. Furthermore, applying a projector method, such as GFMC, to our NQS ansatz further refines the energy to $-0.7296(1)$, demonstrating the high quality of our NQS as a trial wavefunction.
@@ 5.3 Comparison with PEPS
We also benchmark the computational efficiency of NQS against PEPS on identical hardware. The implementation of PEPS can be referred to  (?). As detailed in Figure S2, the expressive power of NQS is similar to fermion PEPS with $D=10$. However, NQS is significantly better than PEPS in terms of computational efficiency under the same hardware condition, as shown in Table S6.
@@ 5.4 Optimizer Ablation Study
To assess the efficacy of our proposed optimizer MARCH, we performed an ablation study comparing its performance against SPRING  (?) and Min-SR  (?). The convergence behavior of both optimizers during the NQS training is presented in Figure S3. These results clearly demonstrate that MARCH achieves significantly faster and more stable convergence compared to SPRING and Min-SR for the systems tested.
@@ 6 Hyperparameter
We present a general settings of NNB training, pretraining, and MARCH training in Table S1, Table S2, and Table S3, respectively. We note that for different systems, the hyperparameters might be slightly different. All code and data will be made openly available upon publication.
@@
Figure S1: (a) Overview of our NQS wavefunction architecture. This figure gives an example of a $2\times 2$ Hubbard model. The color of the sites denotes occupations of electrons on the sites. The process of embedding produces the same number of tokens as the number of sites, which are given as input to the neural network. (b) Details of each building blocks. Transformer will output $K$ sets of backflow orbitals, which are selected according to occupations of electrons. The final wavefunction amplitude is the sum of all determinants.
Figure S2: NQS vs PEPS under the same VMC batchsize. NQS denotes hidden dimension of 256 and 4 determinants, NQS large denotes hidden dimension of 384 and 6 determinants. D denotes bond dimension in PEPS. The expressive power of NQS is similar to PEPS with $D=10$.
Figure S21: Similar as Figure S4 but for $24\times 8$ Hubbard model with $t^{\prime}=-0.2$ under cylinder boundary condition with AFM pinning fields on the edges.
Figure S22: Similar as Figure S4 but for $32\times 6$ Hubbard model with $t^{\prime}=-0.2$ under PBC.
Figure S24: Similar as Figure S4 but for $16\times 10$ Hubbard model with $t^{\prime}=-0.2$ under PBC. We apply temporary pinning field to get vertical stripe.
Systems $8\times 8$ $10\times 10$ $12\times 12$
AFQMC -0.5262(5) -0.5254(3) -0.5246(1)
NQS -0.52582 -0.52492 -0.52440
Table S5: Benchmark energy in pure Hubbard model with OBC. NQS achieve new state-of-the-art results by surpassing fPEPS.
DMRG -0.68537 -0.70550 -0.69840 -0.71250
NQS -0.68325 -0.70307 -0.69952 -0.71309 -0.72212 -0.72747
NQS + GFMC -0.68387(1) -0.70413(1) -0.70055(1) -0.71432(2) -0.72366(2) -0.72885(1)
Table S6: Computation complexity relative to NQS under the same hardward condition in $8\times 8$ lattice. NQS is siginificantly more efficiency than PEPS.
Methods NQS NQS large PEPS D=8 PEPS D=10
Complexity 1×1$\times$1 × 2×2$\times$2 × 7×7$\times$7 × 24×24$\times$24 ×
Table S7: Accuracy on simple systems, where exact solution is possible.
Algorithm  (?)  (?)  (?)  (?)  (?) This work
Lattice $8\times 8$ $4\times 4$ $4\times 4$ $4\times 4$ $8\times 8$ $12\times 12$
Dopping 00 $1/8$ $3/8$ $1/8$ 00 00
Relative Error $0.003$ $0.01$ $0.0005$ $0.001$ $0.005$ $0.0002$
Accuracy $99.7\%$ $99\%$ $99.95\%$ $99.9\%$ $99.5\%$ $99.98\%$

### v2

@@ 2 Pretrain
Instead of randomly initializing the NQS ansatz, we use a pretrain method to provide the transformer wavefunction better initial parameters before optimizing it with MARCH.
We first train a simple neural network similar to neural network backflow (NNB)  (?). A multilayer perceptron (MLP) with two hidden layers maps the occupation number $\mathbf{n}\in\{0,1\}^{2N}$ to a orbital matrix $M_{\text{nnb}}(\mathbf{n})\in\mathbb{R}^{2N\times N_{e}}$, which is restricted to a block-diagonal matrix,
$M_{\text{nnb}}(\mathbf{n})=\begin{pmatrix}M_{\text{nnb}}^{\uparrow}(\mathbf{n})&0\\ 0&M_{\text{nnb}}^{\downarrow}(\mathbf{n})\end{pmatrix},$ (S24)
where the $M_{\text{nnb}}^{\uparrow}(\mathbf{n}),M_{\text{nnb}}^{\downarrow}(\mathbf{n})\in\mathbb{R}^{N\times N_{e}/2}$ are the orbitals of spin-up and spin-down electrons, respectively. Based on the occupation number, $N_{e}$ rows of orbital matrix $M_{\text{nnb}}(\mathbf{n})$ are extracted to form a square orbital matrix $\Phi\in\mathbb{R}^{N_{e}\times N_{e}}$, which is also block-diagonal,
$\Phi=\begin{pmatrix}\Phi^{\uparrow}&0\\ 0&\Phi^{\downarrow}\end{pmatrix}.$ (S25)
The amplitude of NNB ansatz is defined by the determinant of a matrix $\Phi$:
$\psi_{\text{nnb}}(\mathbf{n})=\mathrm{det}[\Phi].$ (S26)
@@
$S_{\text{pre}}=\left\{\mathbf{n}_{i},\ M_{\text{nnb}}(\mathbf{n}_{i})\right\}_{i=0}^{N_{\text{pre}}},$ (S27)
where each entry consists of a configuration $\mathbf{n}_{i}$ and its corresponding orbital matrix $M_{\text{nnb}}(\mathbf{n}_{i})$.
The transformer ansatz is then pretrained in a supervised manner, with the loss function to be the mean square error of the orbital matrix, i.e.,
$\mathcal{L}_{\text{pretrain}}=\frac{1}{N_{\text{pre}}}\sum_{i=1}^{N_{\text{pre}}}\sum_{k=1}^{K}||M^{k}(\mathbf{n}_{i})-M_{\text{nnb}}(\mathbf{n}_{i})||^{2},$ (S28)
with Adam optimizer. We note that in practice, the pretrain dataset can be generated online. Compared to random initialization, a pretrained transformer is much more stable and converges significantly faster.
@@ 3 Pinning Field
To avoid convergence to local minima during optimization, a temporary pinning field is applied. This field takes the form of an antiferromagnetic (AFM) term added to the Hamiltonian $\hat{H}$ on specific columns:
@@
With this method, we can stabilize a vertical or horizontal stripe. From Figure S32 we can see that at the same network size, the vertical stripe is higher in energy than the horizontal stripe, indicating the horizontal stripe indeed is favored on $32\times 8$ lattice with $t^{\prime}=-0.2$ and $\delta=1/8$.
@@ 4.2 Comparison with DMRG
We further compare our NQS method with the DMRG on the $t^{\prime}$ Hubbard model. Specifically, for a cylindrical geometry $16\times 8$ with next-nearest-neighbor hopping $t^{\prime}=-0.2$, our NQS achieves a variational energy of $-0.72791$. This is lower than the DMRG energy of $-0.72747$, obtained with a large bond dimension of $m=40000$.
@@ 4.3 Comparison with PEPS
We also benchmark the computational efficiency of NQS against PEPS on identical hardware. The implementation of PEPS can be referred to Wu et al.  (?). As detailed in Figure S33, the expressive power of NQS is similar to fermion PEPS with $D=10$. However, NQS is significantly better than PEPS in terms of computational efficiency under the same hardware conditions, as shown in Table S3.
@@ 4.4 Optimizer Ablation Study
To assess the efficacy of our proposed optimizer MARCH, we performed an ablation study comparing its performance against SPRING  (?) and Min-SR  (?, ?). The convergence behavior of both optimizers during the NQS training is presented in Figure S34. These results clearly demonstrate that MARCH achieves significantly faster and more stable convergence compared to SPRING and Min-SR for the systems tested.
@@ 5 Hyperparameter
We present general settings of NNB training, pretraining, and MARCH training in Table S5, Table S6, and Table S7, respectively. We note that for different systems, the hyperparameters might be slightly different. All code and data will be made openly available upon publication.
Figure S29: Similar as Figure S11 but for $16\times 10$ Hubbard model with $t^{\prime}=-0.2$ under PBC. We apply a temporary pinning field to get the vertical stripe.
Figure S31: Overview of our NQS wavefunction architecture. (a) This figure gives an example of a $2\times 2$ Hubbard model. The color of the sites denotes the occupations of electrons on the sites. The process of embedding produces the same number of tokens as the number of sites, which are given as input to the neural network. (b) Details of each building block. The transformer will output $K$ sets of backflow orbitals, which are selected according to the occupations of electrons. The final wavefunction amplitude is the sum of all determinants.
Figure S32: Vertical v.s. Horizontal stripe for the $32\times 8$ Hubbard model with $t^{\prime}=-0.2$ and $\delta=1/8$, using NQS with up to $6$ attention heads and determinants.
Figure S33: NQS v.s. PEPS under the same VMC batchsize. NQS denotes the hidden dimension of 256 and 4 determinants, NQS large denotes the hidden dimension of 384 and 6 determinants. D denotes the bond dimension in PEPS. The expressive power of NQS is similar to PEPS with $D=10$.
Systems $8\times 8$ $10\times 10$ $12\times 12$
AFQMC -0.5262(5) -0.5254(3) -0.5246(1)
NQS -0.52582 -0.52492 -0.52440
Table S2: Benchmark energy in pure Hubbard model with OBC. NQS achieve new state-of-the-art results by surpassing PEPS.
DMRG -0.68537 -0.70550 -0.69840 -0.71250
NQS -0.68325 -0.70307 -0.69952 -0.71309 -0.72212 -0.72747
Table S3: Computation complexity relative to NQS under the same hardware condition in $8\times 8$ lattice. NQS is significantly more efficient than PEPS.
Methods NQS NQS large PEPS D=8 PEPS D=10
Wall Time 1×1$\times$1 × 2×2$\times$2 × 7×7$\times$7 × 24×24$\times$24 ×
Table S4: Accuracy on simple systems, where the exact solution is possible.
Algorithm  (?)  (?)  (?)  (?)  (?) This work
Lattice $8\times 8$ $4\times 4$ $4\times 4$ $4\times 4$ $8\times 8$ $12\times 12$
Dopping 00 $1/8$ $3/8$ $1/8$ 00 00
Relative Error $0.003$ $0.01$ $0.0005$ $0.001$ $0.005$ $0.0002$
Accuracy $99.7\%$ $99\%$ $99.95\%$ $99.9\%$ $99.5\%$ $99.98\%$